Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
fiori, derosa, paoletti, tringali, chincarini - 17:31 Tuesday 18 January 2022 (54512) Print this report
Injection coil at NEB, first tests

Today we performed some test injections with the NEB wall coil. The coil has been connected to the amplifier module (the connection is still temporary and some work has to be finalized, for instance the acquistion of monitoring signals and remote switch off are still to be implemented).

We added one magnetic probe (named ENV_NEB_MAG_COIL_N) positioned at the geometric center of the wall coil and oriented along the coil axis, that is along the N arm (Figures 1 and 2). Note that, at first we attached the probe to the SAS separating wall, and the probe was saturating, probabily because of eddy currents induced in aluminum layer of the wall.

Also: we moved magnetometer ENV_NEB_MAG_V onto the tower base, EAST side, first orienting along the N-arm direction and then perpendicular to it (see details in the attached log file). While magnetometers ENV_NEB_MAG_W and ENV_NEB_MAG_N remained in their usual assigned location.

Pictures of the magnetometers are attached.

We injected into the amplifier module (Voltage input) white band passed from 0-1000 Hz (channel ENV_NOISE_NEB).

We rotated the ENV_NEB_MAG_V onto the tower base to the E-W direction and repeated the injection.

DETAILS of the INJECTION are in the attached LOG file

Figure 7 compares transfer functions between the injected noise and the 4 magnetic channels (purple: tower magnetometer oriented North-arm, blue: tower magnetomeer oriented West-arm).

The 1/f slope above 10 Hz or so seems to correspond to the pole of the wall coil. A more precise analisys of data and computation of coil pole based on teoretical and measured R and H parameters has to be done.

Figure 8 shows ffts of the injected magnetic noise.



Images attached to this report
Non-image files attached to this report
Comments to this report:
fiori, Paoletti, Tringali - 17:32 Tuesday 25 January 2022 (54600) Print this report

Today we performed further tests with the NEB coil

  • we investigated a likely reason for the anomalous phase rotation noticed in the measured TF between the ENV_NEB_NOISE and magnetometer signals (see previous entry) (results to be posted).
  • we practiced injecting current noise and voltage noise, at various levels, to get an idea of the maximum colored magnetic noise that we can deal. Both the amplifier saturation limit and the magnetometers saturation limit have to be taken into account. While at high frequency (above 100Hz) injecting in voltage mode allows to avoid amplifier saturation, below 100Hz we soon reach the limit of magnetometers saturation. We find that roughly that in the 0-1000Hz we can inject a noise level that is a factor 100-200 above the quiet ambient noise.
  • we measured the TF between magnetometers in the reference location (N and W magnetometers) and one magnetometers placed in different locations in proximity of the NEB vacuum chamber (data to be analyzed and posted).

Details and gps times of single injections are in the attached log file

Non-image files attached to this comment
Search Help